# A Topological Gaussian Mixture Model for Bone Marrow Morphology in Leukaemia Qiquan Wang1,†, Anna Song1,2, Antoniana Batsivari2, Dominique Bonnet2, and Anthea Monod1

### v1

@@ Appendix B Exploring Alternative Numbers of Clusters and Clustering Algorithms in Texture Decomposition
In Section 4.2, we noted that alternative clustering algorithms, such as GMM and CLARA, and with different numbers of textural clusters, produce similar groupings of the samples into the three phases. In this section, we illustrate the agreement in group allocations with further examples and figures.
@@
In Figures 7 and 9 in Section 4.2, we showed that the texture decomposition with three clusters gave rise to three distinct clusters of samples in the PCA-embedded space corresponding to the different phases. We show with the following Figures 19, 20, and 21, that using different and larger numbers of clusters (10, 15, and 25 respectively) in the analysis nevertheless leads to the same groupings. The phases are colored using green, orange, and red to represent Phases O, I, and II. We observe that the different colored samples can be grouped distinctly on the PCA-embedded plane.
@@ B.2 Alternative Clustering Methods for Texture Decomposition
We now briefly outline the methods. Partition around medoids (PAM), also known as -medoids, is a clustering algorithm that is similar to the -means algorithm with the modification that the centroids now instead can only be points in the dataset. This makes the algorithm more robust to outliers and the computed means are more meaningful, however it is more computationally expensive, especially for large datasets as distance matrices need to be computed for the entire dataset, making the complexity . CLARA is an extension of PAM which deals with the computational complexity of large datasets through subsampling. CLARA subsamples multiple subsets of the original data and within iterations retains the medoids of the subset with the lowest average or total within-cluster dissimilarity. GMM is described in Section 5.1 and the data points are assigned to clusters according to probabilities of belonging to each Gaussian distribution, denoted above by .
@@ B.2.1 Results using GMM Clustering
Using GMM clustering, we obtain results similar to those shown in Figures 22 and 23 for 5 and 10 clusters. For some number of clusters, the CTRL samples can be seen as distinct from the samples of the other two phase groups as in Figure 22(b), while for others, the CTRL samples are clustered with other phase groups as in Figure 23(b). However, from the experimental design, we understand the CTRL samples are distinct from the other samples as they are the unmodified bone samples and so we still maintain a control grouping for these samples.
@@ Appendix C Persistence-Weighted Diagrams
Type | Knee region samples | ||||||
CTRL |
MNC |
U937 |
HL60 |
P1 |
P2 |
Type | Long region samples | ||||||
CTRL |
MNC |
U937 |
HL60 |
P1 |
P2 |
Type | Samples | ||||||
Phase O |
Phase I |
Phase II |
Appendix D UPGMA Trees Based on Distance Between Weighted SDPH Diagrams
In this section, we include the additional UPGMA tree to those found in Section 4.3 constructed from the similarities between different regions of the SDPH diagrams. The tree in Figure 25 is based on similarities between density estimates for the knee region samples, while Figures 26 and 27 are based on similarities in and density estimates. From the clustering derived from the tree, we see that clusters correspond to the phase grouping, as we saw in Section 4.3 for the NW region and we see the distinction between samples of different phases to be much clearer in and compared to .
Appendix E Model Evaluation: Additional Figures and Results
@@
In this section, we provide additional results on our model evaluation discussed in Section 5.2.
E.1 Mixture Size Selection
@@
The number of components of our mixture model is determined by a trade-off between the goodness of fit of the models and the computation time. We explored a range of sizes of models from 2 to 20 components and we cap the maximum size to 20 for generalizability of the models across a range of engraftment levels within the same phase, with larger models being more restrictive. Below, as an example, we present the choice of model size for the SW quadrant. Figure 28 shows the comparison of computation time and model fit measured by the Bayesian information criterion (BIC) with increasing model sizes. We selected BIC as the criterion since it demonstrates greater penalization on the complexity of the model and converges asymptotically to the true model if present. Balancing the need for the model to approximate the data well, and for the algorithm to run in appropriate time with bootstrapping, we found a small model with around 5 components to be optimal. Figure 29 further studies the model fit within each of the phases; we observe “elbows” when the model contains 3 to 5 components, after which the decrease in BIC is noticeably smaller. Transitioning from the Phase 0 model to the Phase I, we also found that incrementing the model size by one led to better model fit.
E.2 Additional Quantitative Evaluation for PH1 NW
In this section, we provide a quantitative evaluation of the coherence of the Phase models to each of the samples for PH1 NW by taking the sum of Hellinger distances or KL-divergence values from each bootstrap subsample to each model. The results are shown in Table 6. Within each column, we see that mostly the rows with the smaller values are those whose samples’ Phases correspond with that of the model with a few exceptions such as for HL60 25% (2). We highlight this correspondence by presenting the values of sums of distances in bold font.
Sample | Phase | Sum of Hellinger distances to model | Sum of KL-divergence to model | ||||
Phase 0 | Phase I | Phase II | Phase 0 | Phase I | Phase II | ||
CTRL_0%(1) | 0 | 4.45 | 8.49 | 14.51 | 1.99 | 6.10 | 21.23 |
CTRL_0%(2) | 0 | 5.46 | 9.70 | 14.64 | 3.32 | 8.38 | 23.69 |
CTRL_0%(3) | 0 | 4.81 | 9.59 | 14.10 | 2.44 | 7.92 | 20.98 |
CTRL_0%(4) | 0 | 5.87 | 8.08 | 16.16 | 3.41 | 5.75 | 28.14 |
U937_1%(1) | I | 8.39 | 6.91 | 14.34 | 6.42 | 4.69 | 21.74 |
U937_1%(2) | I | 8.35 | 6.97 | 15.97 | 6.15 | 4.49 | 28.05 |
U937_7% | I | 11.67 | 5.38 | 15.15 | 12.15 | 2.72 | 21.18 |
U937_8% | I | 10.19 | 5.91 | 15.96 | 8.84 | 2.83 | 25.06 |
U937_10%(1) | I | 11.32 | 5.84 | 15.37 | 11.05 | 2.99 | 22.01 |
U937_10%(2) | I | 10.65 | 6.98 | 13.45 | 10.41 | 4.54 | 17.15 |
U937_10%(3) | I | 10.17 | 5.89 | 14.84 | 9.72 | 3.06 | 21.16 |
HL60_23% | I | 5.86 | 6.67 | 13.96 | 3.17 | 3.91 | 19.54 |
HL60_25%(1) | I | 8.38 | 6.25 | 16.08 | 6.66 | 4.11 | 29.12 |
HL60_25%(2) | I | 15.55 | 10.29 | 20.25 | 26.62 | 13.02 | 54.27 |
P1_10% | I | 7.71 | 5.87 | 13.37 | 4.73 | 2.72 | 16.94 |
P1_40% | II | 8.30 | 8.27 | 11.84 | 5.99 | 5.76 | 13.70 |
P1_44% | II | 14.48 | 13.34 | 10.48 | 17.29 | 13.82 | 10.92 |
P1_51% | II | 16.25 | 16.14 | 9.45 | 18.83 | 18.63 | 7.35 |
P1_60% | II | 10.35 | 11.52 | 9.88 | 7.73 | 10.02 | 8.81 |
P1_76% | I | 9.23 | 7.65 | 14.26 | 8.85 | 6.48 | 24.75 |
P2_59% | II | 16.30 | 17.08 | 9.58 | 19.22 | 21.78 | 8.07 |
P2_88% | II | 14.73 | 13.30 | 10.02 | 16.75 | 13.01 | 9.39 |
P2_90% | II | 13.79 | 15.28 | 9.76 | 14.03 | 17.89 | 8.72 |
MNC_53% | I | 8.04 | 7.64 | 13.42 | 6.41 | 5.87 | 20.57 |
MNC_67% | II | 15.20 | 13.36 | 10.04 | 18.43 | 12.93 | 8.93 |
MNC_75% | II | 14.09 | 16.04 | 9.66 | 14.23 | 19.55 | 7.62 |
MNC_86% | II | 18.22 | 17.20 | 12.42 | 27.23 | 22.19 | 15.55 |
E.3 Additional Figures for Model Evaluation of PH1 NE and PH2 NE Quadrants
We provide additional figures to Section 5.2 on the evaluation of the model on the NE and NE quadrants. Figures 30 and 31 show boxplots of the Hellinger distance between subsamples and phase models, where (a), (b), and (c) show the comparison with the Phase O, I, and II models respectively. We find that with the exception of a few anomalies, samples of each phase most closely resemble its own phase model. This is further supported by Tables 7 and 8.
Sample | Phase | Sum of Hellinger distances to model | Sum of KL-divergence to model | ||||
Phase 0 | Phase I | Phase II | Phase 0 | Phase I | Phase II | ||
CTRL_0%(1) | 0 | 9.26 | 9.25 | 14.96 | 8.49 | 7.59 | 21.68 |
CTRL_0%(2) | 0 | 8.19 | 12.76 | 12.31 | 7.25 | 13.37 | 14.85 |
CTRL_0%(3) | 0 | 9.53 | 13.43 | 12.07 | 8.27 | 14.14 | 13.88 |
CTRL_0%(4) | 0 | 10.80 | 11.29 | 12.91 | 11.06 | 10.40 | 15.49 |
U937_1%(1) | I | 10.15 | 10.16 | 15.08 | 10.08 | 9.27 | 21.59 |
U937_1%(2) | I | 12.31 | 11.13 | 14.69 | 14.13 | 10.42 | 19.85 |
U937_7% | I | 13.45 | 9.10 | 17.53 | 17.52 | 7.72 | 28.95 |
U937_8% | I | 13.45 | 8.84 | 17.50 | 17.16 | 7.13 | 28.70 |
U937_10%(1) | I | 13.99 | 9.56 | 18.57 | 19.30 | 8.82 | 34.00 |
U937_10%(2) | I | 11.08 | 9.59 | 15.68 | 11.52 | 7.78 | 22.32 |
U937_10%(3) | I | 11.41 | 8.97 | 16.13 | 12.08 | 6.81 | 24.27 |
HL60_23% | I | 9.75 | 10.25 | 13.65 | 8.82 | 8.64 | 17.06 |
HL60_25%(1) | I | 12.46 | 9.58 | 17.69 | 14.96 | 8.34 | 30.51 |
HL60_25%(2) | I | 15.93 | 11.82 | 20.41 | 31.39 | 20.91 | 47.47 |
P1_10% | I | 8.27 | 9.37 | 14.96 | 7.02 | 7.88 | 21.83 |
P1_40% | II | 9.44 | 13.71 | 11.68 | 8.53 | 15.24 | 12.81 |
P1_44% | II | 8.42 | 13.55 | 10.59 | 7.13 | 15.08 | 10.82 |
P1_51% | II | 14.19 | 19.06 | 11.14 | 15.79 | 28.54 | 9.68 |
P1_60% | II | 10.72 | 15.21 | 11.22 | 9.78 | 18.01 | 10.52 |
P1_76% | I | 11.16 | 11.50 | 15.67 | 13.46 | 12.88 | 24.71 |
P2_59% | II | 13.51 | 17.99 | 10.38 | 14.80 | 25.66 | 8.89 |
P2_88% | II | 11.14 | 15.82 | 9.83 | 11.36 | 20.46 | 8.80 |
P2_90% | II | 13.58 | 17.28 | 10.43 | 15.91 | 24.47 | 9.31 |
MNC_53% | I | 9.64 | 10.70 | 14.85 | 9.14 | 9.95 | 21.31 |
MNC_67% | II | 10.65 | 14.71 | 11.52 | 9.46 | 17.29 | 11.45 |
MNC_75% | II | 13.96 | 17.53 | 9.95 | 16.14 | 24.66 | 8.29 |
MNC_86% | II | 13.10 | 16.78 | 8.84 | 15.48 | 23.30 | 7.54 |
Sample | Phase | Sum of Hellinger distances to model | Sum of KL-divergence to model | ||||
Phase 0 | Phase I | Phase II | Phase 0 | Phase I | Phase II | ||
CTRL_0%(1) | 0 | 5.70 | 11.51 | 11.81 | 3.11 | 11.77 | 11.83 |
CTRL_0%(2) | 0 | 6.13 | 12.68 | 9.99 | 4.12 | 14.47 | 9.49 |
CTRL_0%(3) | 0 | 7.50 | 11.92 | 9.82 | 5.49 | 11.96 | 8.59 |
CTRL_0%(4) | 0 | 9.24 | 9.52 | 11.02 | 7.42 | 7.56 | 10.01 |
U937_1%(1) | I | 12.56 | 9.48 | 12.83 | 14.13 | 8.37 | 15.69 |
U937_1%(2) | I | 11.92 | 8.87 | 12.97 | 11.73 | 6.83 | 14.87 |
U937_7% | I | 11.66 | 8.76 | 13.81 | 11.15 | 6.36 | 16.01 |
U937_8% | I | 13.52 | 9.49 | 16.07 | 15.51 | 7.80 | 23.12 |
U937_10%(1) | I | 13.72 | 9.76 | 15.11 | 15.38 | 7.96 | 19.54 |
U937_10%(2) | I | 13.05 | 9.87 | 12.94 | 14.33 | 8.13 | 13.91 |
U937_10%(3) | I | 12.98 | 9.05 | 14.03 | 14.04 | 7.01 | 17.26 |
HL60_23% | I | 9.97 | 9.65 | 11.78 | 9.13 | 8.22 | 12.75 |
HL60_25%(1) | I | 11.33 | 8.86 | 13.72 | 11.20 | 6.94 | 16.96 |
HL60_25%(2) | I | 15.14 | 11.00 | 18.18 | 20.58 | 11.18 | 31.46 |
P1_10% | I | 10.08 | 9.11 | 11.23 | 8.67 | 6.84 | 10.57 |
P1_40% | II | 9.13 | 11.38 | 10.15 | 7.85 | 10.86 | 9.33 |
P1_44% | II | 10.93 | 13.20 | 9.12 | 10.36 | 14.83 | 7.22 |
P1_51% | II | 12.47 | 13.96 | 9.78 | 13.17 | 16.67 | 8.07 |
P1_60% | II | 12.96 | 11.67 | 10.70 | 14.36 | 11.34 | 9.47 |
P1_76% | I | 10.94 | 10.08 | 11.33 | 11.18 | 8.94 | 11.48 |
P2_59% | II | 12.72 | 15.68 | 9.78 | 14.51 | 21.39 | 8.63 |
P2_88% | II | 11.78 | 14.14 | 10.09 | 12.95 | 17.60 | 9.52 |
P2_90% | II | 13.19 | 13.96 | 9.90 | 15.76 | 17.15 | 9.50 |
MNC_53% | I | 8.76 | 12.68 | 10.94 | 6.31 | 12.96 | 9.56 |
MNC_67% | II | 12.14 | 12.94 | 10.79 | 13.11 | 14.07 | 9.77 |
MNC_75% | II | 10.92 | 12.25 | 8.88 | 10.30 | 12.72 | 7.98 |
MNC_86% | II | 12.75 | 15.89 | 10.37 | 15.16 | 23.50 | 10.56 |
E.4 Model Evaluation for PH0 SW and PH1 SW Quadrants
The distributions observed in the SW quadrant across the samples were highly similar resulting in the likeness between the fitted GMMs for the three phases. The GMMs are visualized in Figure 32 and the goodness of fit of the fitted model is shown in Figure 33 measured in terms of the Hellinger distance between the fitted model and test subsamples. We find the intervals of the distances overlapping, suggesting little difference between the models.
For the SW quadrant, we observe a similar behavior in that there exists little difference in the patterns and predicted models between the three phases. The model fit and evaluation are shown in Figures 34 and 35.

### v2

@@ Appendix B Exploring Alternative Numbers of Clusters and Clustering Algorithms in Texture Decomposition
In Section 4.2 of the article, we noted that alternative clustering algorithms, such as GMM and CLARA, and with different numbers of textural clusters, produce similar groupings of the samples into the 3 phases. In this section, we illustrate the agreement in group allocations with further examples and figures.
@@
In Figures 7 and 9 in Section 4.2 of the article, we showed that the texture decomposition with 3 clusters gave rise to 3 distinct clusters of samples in the PCA-embedded space corresponding to the different phases. We show with the following Figures 18, 19, and 20, that using different and larger numbers of clusters (10, 15, and 25 respectively) in the analysis nevertheless leads to the same groupings. The phases are colored using green, orange, and red to represent Phases O, I, and II. We observe that the different colored samples can be grouped distinctly on the PCA-embedded plane.
@@ B.2 Alternative Clustering Methods for Texture Decomposition
We now briefly outline the methods. Partition around medoids (PAM), also known as -medoids, is a clustering algorithm that is similar to the -means algorithm with the modification that the centroids now instead can only be points in the dataset. This makes the algorithm more robust to outliers and the computed means are more meaningful, however it is more computationally expensive, especially for large datasets as distance matrices need to be computed for the entire dataset, making the complexity . CLARA is an extension of PAM which deals with the computational complexity of large datasets through subsampling. CLARA subsamples multiple subsets of the original data and within iterations retains the medoids of the subset with the lowest average or total within-cluster dissimilarity. GMM is described in Section 5.1 of the article and the data points are assigned to clusters according to probabilities of belonging to each Gaussian distribution, denoted above by .
@@ B.2.1 Results using GMM Clustering
Using GMM clustering, we obtain results similar to those shown in Figures 21 and 22 for 5 and 10 clusters. For some number of clusters, the CTRL samples can be seen as distinct from the samples of the other two phase groups as in Figure 21(b), while for others, the CTRL samples are clustered with other phase groups as in Figure 22(b). However, from the experimental design, we understand the CTRL samples are distinct from the other samples as they are the unmodified bone samples and so we still maintain a control grouping for these samples.
@@ Appendix C Persistence-Weighted Diagrams
-1cm
Type
Knee region samples
CTRL
MNC
U937
HL60
P1
P2
-1cm
Type
Long region samples
CTRL
MNC
U937
HL60
P1
P2
-1cm
Type
Samples
Phase O
Phase I
Phase II
Appendix D Consensus Hierarchical Clustering
In this section, we present the additional consensus hierarchical trees for PH2 KDEs. Figure 24 shows the consensus hierarchical clusters obtained across bootstrap resamplings, while Figure 25 depicts consensus across different linkage methods. In both cases, we extract the 4 clusters by cutting the dendrograms at appropriate heights. Among the resultant clusters, 2 clusters correspond exclusively to Phase II and Phase I samples, respectively, with the P1 40% sample once again appearing outside the sole Phase II group as an outlier. The remaining 2 clusters are composed primarily of Phase O and Phase I samples, reflecting the agreement to the Phase labels.
Appendix E Continuous Modeling of Disease Progression
While our primary analyses rely on 3 inferred discrete disease phases, we recognize that AML progression is inherently continuous. A regression-based framework could, in principle, offer a more fine-grained representation of disease state. However, several biological and experimental factors present nontrivial challenges to implementing such an approach robustly in our setting. Firstly, each data sample in our study corresponds to a single terminally sacrificed mouse, eliminating the possibility of within-subject longitudinal modeling. Moreover, the biological heterogeneity across miceāincluding differences in immune response, genetic background, and AML subtypeācomplicates the assumption of a shared continuous timeline across animals. Furthermore, engraftment levels, often used as a proxy for disease burden, were measured from skeletal sites distinct from the imaged femur and are thus subject to both biological and spatial noise. These values should be interpreted as imprecise, interval-censored indicators of progression, without clearly defined temporal bounds.
Despite these limitations, we investigated a prototype continuous-time model based on a spline-interpolated time-varying Gaussian Mixture Model (GMM). In this framework, each time point is associated with a density
where , , and are the mixture weights, means, and covariances, respectively, and the weights and means are then smoothed using spline functions for interpolation. While this provides a flexible framework for capturing smooth temporal evolution in the data, it is also affected by the volatility of the estimated Gaussian parameters and by the limited number of temporal time points, which can undermine stability and reliability in downstream inference.
We evaluated predictive performance using leave-one-out cross-validation (LOOCV) and the results are presented in Figure 26. Although the model exhibited some alignment with the engraftment levels, its overall accuracy was limited. Performance was affected by several factors: the sparsity and repetition of time points complicated spline fitting; numerical instability arose during inversion of the spline mappings for prediction; and noise in the engraftment labels further impeded reliable inference. More sophisticated strategiesāsuch as incorporating stochastic interpolation, treating time as a latent variable in a Bayesian framework, or modeling measurement error explicitlyāmay offer improvements. However, these approaches substantially increase model complexity and risk overfitting given the limited sample size. For these reasons, we adopt a more interpretable and statistically stable discretization into 3 phases, which preserves meaningful biological structure without overburdening the data.
Appendix F Model Interpretation: Additional Figures and Results
@@
In this section, we provide additional results on our model evaluation discussed in Section 5.2 of the article.
F.1 Mixture Size Selection
@@
The number of components of our mixture model is determined by a trade-off between the goodness of fit of the models and the computation time. We explored a range of sizes of models from 2 to 20 components and we cap the maximum size to 20 for generalizability of the models across a range of engraftment levels within the same phase, with larger models being more restrictive. Below, as an example, we present the choice of model size for the SW quadrant. Figure 27 shows the comparison of computation time and model fit measured by the Bayesian information criterion (BIC) with increasing model sizes. We selected BIC as the criterion since it demonstrates greater penalization on the complexity of the model and converges asymptotically to the true model if present. Balancing the need for the model to approximate the data well, and for the algorithm to run in appropriate time with bootstrapping, we found a small model with around 5 components to be optimal. Figure 28 further studies the model fit within each of the phases; we observe āelbowsā when the model contains 3 to 5 components, after which the decrease in BIC is noticeably smaller. Transitioning from the Phase O model to the Phase I, we also found that incrementing the model size by one led to better model fit.
F.2 Model Evaluation for PH0 SW and PH1 SW Quadrants
In this section, we show our model fitted to the PH0 SW and PH1 SW quadrants. In the PH0 SW quadrant, the axes correspond to values of and , representing vessel thickness at bumps and narrowings along the vessel. Across the Phase O, I, and II models as shown in Figure 29, we observe relatively stable distributions with minimal shift in density across the phases. This suggests that the morphological features captured in this quadrant remain largely unaffected during AML progression.
Within the PH1 SW quadrant, the axes correspond to the values of and , quantifying vessel thicknesses at narrowings and dimples, we observe a subtle but consistent trend. As shown in Figure 30, compared to Phase O, the Phase I and Phase II models exhibit a slight shift of the high-probability region toward lower values of both and . This suggests a modest reduction in vessel size at these critical points, indicating that as AML progresses, vessels may become more constricted or exhibit more prominent surface depressions in the vessel structure.